What is the answer and how was it figured out for this 3x - 17 = 7x - 9
Thank You,
Gerry

Answer by Abbey(339)

the goal is to get something that looks like x=(some number), so subtract 3x from both sides of the equation:
-17=7x-3x-9
-17=4x-9
then add 9 to both sides:
-17+9=4x
-8=4x
then divide both sides by 4
-2=x
so the solution is
x=-2